# Style Content Decomposition-based Data Augmentation for Domain Generalizable Medical Image Segmentation

### v2

@@ Abstract
Due to domain shifts across diverse medical imaging modalities, learned segmentation models often suffer significant performance degradation during deployment. These domain shifts, typically caused by variations in imaging systems, generally comprise two principal components: 1) ”style” shifts, referring to global disparities in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local discrepancies in anatomical structures. To address domain shifts in medical image segmentation, a core challenge arises: how can we decouple the factors within images that determine their ”style” and ”content” components? To this end, we first propose a linear style-content decomposition method that factorizes an image into style codes and content maps, explicitly modeling the ”style” and ”content” components. Building on this, we introduce a Style-Content decomposition-based data augmentation algorithm (StyCona), which leverages this decomposition strategy to guide augmentation of both the global style and local content of source-domain images, enabling the training of a well-generalized model for domain-generalizable medical image segmentation. StyCona is a simple yet effective plug-and-play module that substantially improves model generalization without requiring additional training parameters or modifications to segmentation model architectures. Experiments on cardiac magnetic resonance imaging and fundus photography segmentation tasks, with single and multiple target domains respectively, demonstrate the effectiveness of StyCona and its superiority over state-of-the-art domain generalization methods. The code will be released at https://github.com/Senyh/StyCona.
Introduction
@@
Medical image segmentation is critical for computer-aided diagnosis. Driven by large and diverse annotated datasets, deep learning-based segmentation models have achieved remarkable progress (Ronneberger, Fischer, and Brox 2015; Milletari, Navab, and Ahmadi 2016). However, data distribution mismatches (a.k.a., domain shifts), caused by diverse imaging protocols, equipment vendors, or patient populations, etc., between source (training) and target (testing) domains hinder the generalizability of trained models for clinical deployment (Castro, Walker, and Glocker 2020; Zhou et al. 2022; Guan and Liu 2021).
@@
Figure 1: Illustration of style codes and content maps for bSSFP and LGE MRI sequences: a) original images showing disparities in both style (global image appearance) and content (local anatomical structures), b) image histograms reflecting domain shifts in terms of pixel intensities, c) style-swapped images generated by swapping the two original images’ style codes (top: bSSFP image with LGE style; bottom: LGE image with bSSFP style), d) content maps of the original images, and e) t-SNE (Van der Maaten and Hinton 2008) visualization of domain shifts (top: style shift; bottom: content shift).
@@
Domain generalization (DG) has emerged as a promising strategy to mitigate domain shifts by training models solely on source-domain data and generalizing them to unseen domains. Current domain generalization (DG) methods encourage segmentation models to learn domain-invariant representations through representation learning (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gao et al. 2022, 2023), data augmentation (a.k.a., domain randomization) (Ouyang et al. 2022; Zhou et al. 2021; Chen et al. 2023, 2025), as well as the combination of both schemes (Gu et al. 2023; Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023). Specifically, the first line of work focuses on explicitly learning domain-invariant information via deterministic (Chen et al. 2019, 2020; Pei et al. 2021) or statistical (Gao et al. 2022, 2023) modeling. While these methods yield intuitive decoupled features on source-domain data, domain shifts pose a dilemma: the disentanglement models, constrained by domain-specific training, often fail to generalize their disentanglement capabilities across different domains. In contrast, data augmentation-based DG methods can circumvent this issue by expanding the source-domain data distribution and thus implicitly encouraging models to capture domain-invariant information. These methods leverage techniques such as random convolution (Xu et al. 2021b; Ouyang et al. 2022; Choi et al. 2023), and feature statistics editing (Zhou et al. 2021; Chen et al. 2023, 2025), etc. The third line of research employs learning-based (Gu et al. 2023) or rule-based (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023) disentanglement to decompose images into domain-invariant contents and domain-specific styles, followed by style augmentation based on the disentangled components. However, the former depends on additional encoders and decoders, as well as access to multiple domain images and their corresponding labels. These requirements are computationally intensive and may be infeasible in practice, e.g., single-source domain generalization or domain-label unavailable scenarios. In contrast, the latter assumes invariance in the Fourier phase spectrum, which limits its effectiveness for data augmentation to alleviate domain shifts. Moreover, the specific components contributing to domain shifts remain underexplored. This raises a question: Can a deeper understanding of the underlying nature of domain shifts guide the design of more effective data augmentation methods for domain generalizable medical image segmentation?
@@
To answer this question, we begin with a visual investigation of domain shifts in medical images. As shown in Fig.1(a-b), domain shifts generally manifest as variations in pixel intensity, comprising two major components: 1) style shift, referring to global disparities in image characteristics such as brightness, contrast, and color; and 2) content shift, referring to local anatomical discrepancies caused by differences in tissue visualization across imaging systems. Inspired by these observations, rather than learning-based (Gu et al. 2023) or rule-based (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023) disentanglement, we explicitly model the ”style” and ”content” components of an image by linearly decomposing it into a style code and content maps. We reveal that 1) the style code captures global image characteristics within a given anatomical structure and 2) the content maps describe the image’s anatomical structure. To provide a clearer illustration, we visualize style-swapped images (by exchanging two images’ style codes) and content maps in Fig. 1(c-d); it can be observed that the content maps delineate the anatomical structures of the original bSSFP and LGE images, and the style-swapped images transfer appearance between each other while preserving the original anatomical structures. Furthermore, the shift observed in the style codes is more significant than that in the content maps [Fig. 1(e)], as global variations in image characteristics are generally more substantial than local differences in anatomical structures. These results align with the aforementioned intuition. Based on this, we categorize domain shifts into two components: 1) style shifts (global image property variations) as indicated by deviations in the style codes, and 2) content shifts (local anatomical structure discrepancies) as reflected in differences between the content maps. Given that explicitly measuring and reducing these two types of shifts is infeasible due to the unavailability of target domain data during training, we instead aim to perturb the ”style” and ”content” components of source domain images for generating images from diverse domains (simulating patients undergoing different imaging systems) to train a domain generalizable medical image segmentation model.
@@
To this end, we propose a style content decomposition-based data augmentation algorithm (StyCona) to advance domain generalizable medical image segmentation. Specifically, StyCona performs perturbations on the ”style” (global image characteristics) and ”content” (local anatomical structures) by 1) blending style codes and 2) mixing content maps. We evaluate StyCona on cross-domain cardiac magnetic resonance imaging (MRI) segmentation and optic cup (OC)/optic disk (OD) fundus photography segmentation tasks. Experimental results demonstrate that StyCona is a promising solution for domain generalizable medical image segmentation and achieves state-of-the-art performance.
@@
    To the best of our knowledge, this is the first work that decomposes domain shifts into style shifts and content shifts through linear decomposition of images in domain generalizable medical image segmentation tasks.
@@
    We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. Our method can be easily integrated into off-the-shelf medical image segmentation backbones and effectively mitigates the two types of domain shifts.
@@
Related work
Domain Generalization
Domain generalization (DG) aims to train a model on a source domain and enable it to generalize to unseen target domains. Numerous DG methods have been proposed for both natural and medical image scenarios (Zhou et al. 2022; Castro, Walker, and Glocker 2020). Specifically, these methods enforce the learning of domain-invariant representations through various approaches, including: deterministic or statistical modeling of domain-invariant and domain-specific features (Gu et al. 2023; Gao et al. 2022, 2023); frequency transformation-based mix-up, which combines Fourier amplitude spectra from different images for style augmentation (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024); random convolution-based data augmentation, which perturbs pixel intensities via convolution neural networks with random parameters (Xu et al. 2021b; Ouyang et al. 2022; Choi et al. 2023); as well as feature statistics editing-based style transfer, which perturbs image- or feature-level statistics (e.g., mean and standard deviation) to achieve style or content augmentation (Zhou et al. 2021; Chen et al. 2023, 2025). These techniques are generally applicable to both image recognition and semantic segmentation tasks, with the latter requiring additional consideration for transformation equivalence. However, the fundamental nature of domain shifts, crucial for guiding the development of more effective DG approaches, remains underexplored. In this paper, we propose a novel framework that first decomposes domain shifts into style and content shifts through linear modeling of images and then designs a novel data augmentation algorithm to mitigate both types of domain shifts.
Content-Style Decomposition
Content-style decomposition disentangles features into domain-invariant content and domain-specific style components for addressing the domain shift issue. In the context of semantic segmentation, this technique has been widely employed in the design of both unsupervised domain adaptation (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Yang and Soatto 2020) and domain generalization (Gu et al. 2023; Gao et al. 2022, 2023; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024) frameworks. Research in this area generally falls into two categories: 1) learning-based (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gu et al. 2023; Gao et al. 2022, 2023) and 2) rule-based approaches (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024). Learning-based methods rely on additional deep neural network-based encoders and decoders, with specific constraints (e.g., contrastive learning on decoupled components (Gu et al. 2023)), to achieve feature disentanglement either deterministically (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gu et al. 2023) or statistically (Gao et al. 2022, 2023). Rule-based methods adopt the Fourier transformation to perform direct decomposition. However, they impose the constraint that phase spectra should be fixed to ensure the invariance of the original contents during reconstruction. This constraint limits their flexibility and effectiveness for data augmentation. In contrast, we propose a simple yet effective content-style decomposition method that linearly factorizes an image into a style code and content maps, enabling style-content augmentation through perturbations on the decoupled components.
Figure 2: Schematic diagram of the proposed style content decomposition-based data augmentation algorithm (StyCona). StyCona includes three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation. $x_{i}$ and $x_{j}$ represent original images, and $\bar{\tilde{x}}$ denotes a StyCona-augmented sample.
Methodology
@@
Overview
We assume that domain shifts in medical images resulting from variations in image acquisition processes generally comprise two principal components: 1) ”style” shifts, referring to global variations in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local changes in image characteristics and are usually associated with differences in the visibility of specific tissues. Under this assumption, a key problem arises for addressing the domain shifts: how can we decouple the factors within images that determine the ”style” and ”content” components? Although several approaches have been proposed to address this problem, they either rely on auxiliary disentanglement networks (Gu et al. 2023; Gao et al. 2022, 2023) or impose constraints on the perturbation of the decoupled components (Xu et al. 2021a). Motivated by the effectiveness of linear decomposition in image compression (Andrews and Patterson 2003; Ranade, Mahabalarao, and Kale 2007) and restoration (Zhu et al. 2014; Guo et al. 2015; Gao and Zhuang 2020), we propose to linearly decompose an image into its style code and content maps, from which we validate both aforementioned types of domain shifts, as shown in Fig. 1. Based on the linear decomposition, we introduce a style content decomposition-based data augmentation algorithm, termed StyCona, for domain generalizable medical image segmentation. The training process for leveraging StyCona to train a domain-generalizable medical image segmentation model is summarized in Algorithm 1.
Style Content Decomposition-based Data Augmentation (StyCona)
@@
The core idea behind StyCona is that the style and content shifts can be mitigated through style and content augmentations, respectively. To this end, the StyCona algorithm mainly includes two components: 1) image decomposition and 2) style content augmentation. Specifically, the image decomposition is achieved using singular value decomposition (SVD) (Klema and Laub 1980) to factorize an image into its singular values (style codes) and rank-one matrices (content maps) The style content augmentation is then performed by perturbing style codes and content maps, respectively. As illustrated in Fig. 2, we realize the StyCona algorithm with three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation.
Singular Value Decomposition
For an image $x$, its SVD is formulated as:
$x=\sum_{k}\underbrace{\sigma^{k}}_{style\,code}\,\,\underbrace{\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}}_{content\,map}$ (1)
where each singular value $\sigma^{k}$ acts as a style code and the rank-one matrix $\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}$ serves as a content map for the subsequent style and content augmentation.
Algorithm 1 StyCona for domain generalizable medical image segmentation
Input: $\mathcal{D}=\{(x_{i},y_{i})_{i=1}^{N}\}$
Parameter: $\theta$ for the segmentation model $f(\cdot;\theta)$
Output: Trained model $f(\cdot;\theta)$
1:for each iteration do
2:  # 1) Extract style codes and content maps from current batch original image $x_{i}$ and auxiliary image $x_{j}$ ($i\neq j$) by Eq. (1)
3:  $\boldsymbol{\sigma}_{i}\leftarrow x_{i}\,\,and\,\,\{\boldsymbol{u}^{r}_{i}\times\boldsymbol{v}^{r}_{i}\}\leftarrow x_{i}$
4:  $\boldsymbol{\sigma}_{j}\leftarrow x_{j}\,\,and\,\,\{\boldsymbol{u}^{r}_{j}\times\boldsymbol{v}^{r}_{j}\}\leftarrow x_{i}$
5:  # 2) Style Augmentation by Eq. (2)
6:  $\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$
7:  # 3) Content Augmentation by Eq. (3)
8:  $\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$
9:  # 4) Obtain StyCona images via inverse singular value decomposition
10:  $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$
11:  # 5) Forward propagation to obtain the predictive label for the StyCona image $\bar{\tilde{x}}_{i}$
12:  $\hat{\bar{\tilde{y}}}_{i}\leftarrow f(\bar{\tilde{x}}_{i},\theta)$
13:  # 6) Loss supervision by Eq. (4)
14:  Back-Propagate $\mathcal{L}$
15:  Update $\theta$
16:  return $f(\cdot;\theta)$
17:end for
Style Augmentation
@@
Style augmentation aims to alleviate style shifts (global variations in image properties). As depicted in Fig. 2, we perform style augmentation on an image $x_{i}$ by blending its style codes (i.e., singular values) with those of an auxiliary image $x_{j}$ in an element-wise manner:
$\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$ (2)
@@
where the weight $\alpha\sim U(0,1)$ controls the mixing strength, $\sigma_{i}^{k}$ represents the $k_{th}$ singular value of image $x_{i}$, $i/j$ indicates the sample index and $i\neq j$. After that, a style-augmented image $\tilde{x}_{i}$ can be generated using the perturbed singular values and the original singular vectors by Eq. (1).
Content Augmentation

### v3

@@ Abstract
Due to domain shifts across diverse medical imaging modalities, learned segmentation models often suffer significant performance degradation during deployment. We posit that these domain shifts can generally be categorized into two main components: 1) ”style” shifts, referring to global disparities in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local discrepancies in anatomical structures. To address the domain shifts in medical image segmentation, we first factorize an image into style codes and content maps, explicitly modeling the ”style” and ”content” components. Building on this, we introduce a Style-Content decomposition-based data augmentation algorithm (StyCona), which performs augmentation on both the global style and local content of source-domain images, enabling the training of a well-generalized model for domain generalizable medical image segmentation. StyCona is a simple yet effective plug-and-play module that substantially improves model generalization without requiring additional training parameters or modifications to segmentation model architectures. Experiments on cardiac magnetic resonance imaging and fundus photography segmentation tasks, with single and multiple target domains respectively, demonstrate the effectiveness of StyCona and its superiority over state-of-the-art domain generalization methods. The code is available at https://github.com/Senyh/StyCona.
keywords:
Medical Image Segmentation, Domain Generalization, Style Code, Content Map
1 Introduction
@@
Medical image segmentation is critical for computer-aided diagnosis. Driven by large and diverse annotated datasets, deep learning-based segmentation models have achieved remarkable progress (ronneberger2015u; milletari2016v). However, data distribution mismatches (a.k.a., domain shifts), caused by diverse imaging protocols, equipment vendors, or patient populations, etc., between source (training) and target (testing) domains hinder the generalizability of trained models for clinical deployment (castro2020causality; zhou2022domain; guan2021domain).
@@
Figure 1: Illustration of style codes and content maps for bSSFP and LGE MRI sequences: a) original images showing disparities in both style (global image appearance) and content (local anatomical structures), b) image histograms reflecting domain shifts in terms of pixel intensities, c) style-swapped images generated by swapping the two original images’ style codes (top: bSSFP image with LGE style; bottom: LGE image with bSSFP style), d) content maps of the original images, and e) t-SNE (van2008visualizing) visualization of domain shifts (top: style shift; bottom: content shift).
@@
Domain generalization (DG) aims to mitigate domain shifts by training models on source-domain data and generalizing them to unseen target domains. Existing DG studies generally promote the learning of domain-invariant representations through two major paradigms: representation learning (chen2019synergistic; chen2020unsupervised; pei2021disentangle; peng2022unsupervised; gao2022joint; gao2023bayeseg) and data augmentation (a.k.a., domain randomization) (ouyang2022causality; zhou2021mixstyle; chen2023treasure; chen2025constyx; gu2023cddsa; yang2020fda; xu2021fourier; li2023frequency). The former explicitly explores domain-invariant information via deterministic (chen2019synergistic; chen2020unsupervised; pei2021disentangle) or statistical (gao2022joint; gao2023bayeseg) modeling. Although such methods yield intuitive decoupled features on source-domain data, their reliance on domain-specific training often hinders the generalization of disentanglement capabilities across different domains. In contrast, data augmentation-based DG methods alleviate this limitation by expanding the source-domain data distribution and thus implicitly encouraging models to excavate domain-invariant features. As a result, this paradigm has become the mainstream approach in the field. Specifically, these methods typically employ techniques such as Fourier transformation (yang2020fda; xu2021fourier; li2023frequency), random convolution (xu2021robust; ouyang2022causality; choi2023progressive), and feature statistics editing (zhou2021mixstyle; chen2023treasure; chen2025constyx), etc to achieve style augmentation on source-domain data. However, the underlying components involved in medical images remain underexplored. Modeling these components can provide valuable insights into the nature of domain shifts and guide the design of more effective data augmentation strategies.
@@
To bridge this gap, we introduce a style–content decomposition strategy that decomposes an image into style codes and content maps, revealing that 1) the style code captures global image characteristics within a given anatomical structure and 2) the content maps describe the image’s anatomical structure. Fig. 1(c-d) visualizes style-swapped images (obtained by exchanging the style codes between two images) and content maps. It can be observed that the content maps delineate the anatomical structures of the original bSSFP and LGE images, while the style-swapped images exchange appearance between domains but preserve the original anatomical structures. Furthermore, the shift observed in the style codes is more significant than that in the content maps [Fig. 1(e)], as global variations in image characteristics are generally greater than local differences in anatomical structures. Based on these observations, we categorize domain shifts in medical images into two components: 1) style shifts (global image property variations) as indicated by deviations in the style codes, and 2) content shifts (local anatomical structure discrepancies) as reflected in differences between the content maps. Since quantitatively measuring and reducing these two types of shifts is infeasible without access to target domain data during training, we instead perturb the ”style” and ”content” components of source domain images to generate augmented images from diverse domains (simulating patients undergoing different imaging systems), enabling the training of a well-generalized medical image segmentation model.
@@
To this end, we propose a style content decomposition-based data augmentation algorithm (StyCona) to advance domain generalizable medical image segmentation. Specifically, StyCona performs perturbations on the ”style” (global image characteristics) and ”content” (local anatomical structures) of an image by 1) blending style codes and 2) mixing content maps, respectively. StyCona generates augmented images with diverse styles and contents while preserving their semantic information, enabling the simulation of images from a wide range of domains for training a well-generalized segmentation model. We evaluate StyCona on cross-domain cardiac magnetic resonance imaging (MRI) segmentation and optic cup (OC)/optic disk (OD) fundus photography segmentation tasks. Experimental results demonstrate that StyCona is a promising solution for domain generalizable medical image segmentation.
@@
    We propose a style–content decomposition strategy to model the underlying components of medical images and provide a deep insight into domain shifts.
@@
    We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. StyCona can be easily integrated into off-the-shelf medical image segmentation backbones and effectively mitigates domain shifts.
@@
Figure 2: Schematic diagram of the proposed style content decomposition-based data augmentation algorithm (StyCona). StyCona includes three steps: 1) style content decomposition, 2) style augmentation, and 3) content augmentation. $x_{i}$ and $x_{j}$ represent original images, and $\bar{\tilde{x}}$ denotes a StyCona-augmented sample.
2 Methodology
@@
2.1 Style Content Decomposition-based Data Augmentation (StyCona)
@@
Domain shifts in medical images can be characterized two principal components: 1) ”style” shifts, referring to global variations in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local changes in image characteristics and are usually associated with differences in the visibility of specific tissues. Based on this assumption, we propose StyCona to mitigate both style and content shifts through corresponding style and content augmentations. As illustrated in Fig. 2, StyCona mainly includes three steps: 1) style content decomposition, 2) style augmentations, and 3) content augmentation. Specifically, the style-content decomposition is performed using singular value decomposition (SVD) (klema1980singular), which factorizes an image into its singular values (style codes) and rank-one matrices (content maps). Then, the style and content augmentations are achieved by perturbing the style codes and content maps, respectively.
2.1.1 Style Content Decomposition
Low-rank matrices of an image correspond to its principal components abdi2010principal; candes2011robust. Based on this idea, prior work has employed rank-one matrices as the principal parts of an image for restoration tasks gao2020rank. We leverage SVD to show that an image’s singular values encode its style codes, while the associated rank-one matrices represent its content maps:
$x=\sum^{R}_{r}\underbrace{\sigma^{r}}_{style\,code}\,\,\underbrace{\boldsymbol{u}^{r}{\boldsymbol{v}^{r}}^{\mathsf{T}}}_{content\,map}$ (1)
where $R$ denotes the rank of the image, corresponding to the number of non-zero singular values. SVD provides the optimal rank-one decomposition in the least-squares sense. It guarantees that the set $\boldsymbol{\Phi}=\{\boldsymbol{u}^{1}{\boldsymbol{v}^{1}}^{\mathsf{T}},\boldsymbol{u}^{2}{\boldsymbol{v}^{2}}^{\mathsf{T}},...,\boldsymbol{u}^{R}{\boldsymbol{v}^{R}}^{\mathsf{T}}\}$ forms a complete basis for the column and row spaces of $x$.
  • Content Map. Each rank-one matrix in $\boldsymbol{\Phi}$ functions as a basis pattern of the image, and any image structure can be represented as a linear combination of these patterns. Thus, $\boldsymbol{\Phi}$ completely defines the anatomical structures of the image.
  • Style Code. Each scalar $\sigma^{r}$ acts as a global multiplier for the corresponding basis pattern $(\boldsymbol{u}^{r}{\boldsymbol{v}^{r}}^{\mathsf{T}})$. It modulates the magnitude of pixel intensity in the image, scaling the values of the basis pattern proportionally.
As illustrated in Fig. 1, swapping the style codes of the bSSFP and LGE images results in style transfer between them, while summing all content maps of one of the images reconstructs its complete anatomical structure.
2.1.2 Style Augmentation
@@
Style augmentation aims to alleviate style shifts (global variations in image properties). As depicted in Fig. 2, we perform style augmentation on an image $x_{i}$ by blending its style codes with those of an auxiliary image $x_{j}$ in an element-wise manner:
$\tilde{\sigma}_{i}^{r}=\alpha\times\sigma_{i}^{r}+(1-\alpha)\times\sigma_{j}^{r}$ (2)
@@
where the weight $\alpha\sim U(0,1)$ controls the mix strength, $\sigma_{i}^{r}$ represents the $r_{th}$ singular value of image $x_{i}$, and $i/j$ indicates the sample index ($i\neq j$).
2.1.3 Content Augmentation